A balloon at the top of Mount Logan occupies a volume of 775mL at a temperature of -28°C and a pressure of 92.5kPa. What is the pressure at the bottom of the mountain if the same balloon has a volume of 825mL at a temperature of 15°C? 


1
Expert's answer
2021-01-26T03:23:26-0500

-28oC = 273 - 28 = 245 K

15oC = 273 + 15 = 288 K

According to the combined gas law,


"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}"


Therefore,


"p_2=\frac{p_1V_1T_2}{T_1V_2}=\frac{92.5kPa\times775mL\times288K}{245K\times825mL}=102kPa"


Answer: 102 kPa